# Spatially flat cosmological quantum spacetimes

### v1

@@ 3.2 Explicit form of the $SO(3,1)$ generators
It was shown in [12] that in the minimal representation $n=0$, the generators of rotations $M^{ij}$ are given by [^4] Note that the generators $T^{\mu}$ are no longer hermitian for $n=0$ because $R=\textup{i}r$ in this case. The $SO(3,1)$ generators, however, remain hermitian. $T^{\mu}$ can be made hermitian for $n=0$ via the redefinition $T^{\mu}=\frac{1}{r}(M^{\mu 0}+M^{\mu 4})$, and adjusting the formulas for the $SO(3,1)$ generators accordingly.
@@
where $c$, $d$ and $f$ are functions of the $SO(3)$-invariant quantities $Y^{0}$, $T^{0}$ and $Y^{k}Y_{k}$. However, their semi-classical limits $m^{ij}\sim M^{ij}$ for $n\gg 0$ and also in the minimal case $n=0$ can be determined explicitly. In Appendix B, we derive
@@
A simple explicit realization of the minimal $n=0$ representation in terms of three canonical generators is given in [18].
@@ 3.3 Fiber projection
Fiber projections of polynomials of higher order in $t^{\mu}$ can be computed using Wick type formulas as in [34, 31].
@@ 3.4 Derivatives
In this subsection, we will construct suitable derivative operators acting on the algebra of $\mathfrak{hs}$-valued functions on spacetime, analogous to the construction for k=1k=-1 in [34]. We will always assume $n\gg 0$. To obtain more transparent formulas, it is sometimes convenient to use the variables $x^{a}$ suitable to describe the fuzzy hyperboloid and k=1k=-1 spacetimes instead of $y^{\mu}$. However, in all formulas of the following considerations, $x^{a}$ should be understood as $x^{a}(y^{\mu})$ via the relations
@@
Moreover, Eq. (3.24) and the Poisson brackets (A.4) yield
@@ 3.5 Reconstruction of divergence-free vector fields
where $\eth_{a}$ is the tangential derivative on $H^{4}$ defined in (2.7); and that conversely any vector field $V^{\mu}$ on ${\mathcal{M}}^{1,3}_{k=-1}$ possesses a lift to $H^{4}$ so that (3.30) is satisfied [34]. Note that the vector fields may be $\mathfrak{hs}$ valued. Therefore, there is a correspondence between tangential and divergence-free vector fields on $H^{4}$ to divergence-free vector fields on ${\mathcal{M}}^{1,3}_{k=-1}$. In the following, we establish the analogous correspondence for the $k=0$ case.
@@ 4.5 Local normal coordinates
Due to the $\mathfrak{hs}$ components, the physical meaning of a metric or frame is not clear a priori. As shown in [34], it is always possible to locally eliminate these $\mathfrak{hs}$ components, in local normal coordinates. We illustrate this in the following for the present backgrounds.
@@ 4.7 Physical scales and consistency conditions
On the non-commutative level, the leading contribution to $\lvert[Y^{\mu},Y^{\nu}][Y_{\mu},Y_{\nu}]\rvert$ is the mixed spatial-temporal entry [34]. This should remain true on the semi-classical level, which is easy to verify if the ratio $\tfrac{\alpha}{\beta}$ does not grow too fast. In this case, we generically find
@@
Finally, we require the dilaton to be either constant or increasing with $y^{0}$, corresponding to weak Yang-Mills coupling $g_{\rm YM}\sim\rho^{-1}$ [35]. Using the formula (4.13), this means that $3(1+\mu)+\lambda\geq 0$. This condition might not be strictly necessary, but we will find later that it only excludes some shrinking FLRW spacetimes, which are anyway not interesting from a physical point of view.
@@ 5 FLRW quantum spacetimes and the IKKT model
Let us finally describe how the cosmological quantum spacetimes described above could be incorporated into the IKKT model [^6] It is sometimes questioned whether the IKKT mode admits any non-trivial backgrounds or saddle points with reduced dimension. These doubts are resolved by the example of flat $D3$ BPS branes realized by the Moyal-Weyl quantum plane ${\mathbb{R}}^{3,1}_{\theta}$, leading to noncommutative ${\mathcal{N}}=4$ SYM [1].. To this end, consider the bosonic part of the IKKT action (1.5),
@@
Since $\Box_{\Phi}=[\Phi^{I},[\Phi_{I},\cdot]]$ is a positive operator, non-trivial solutions for the $\Phi^{I}$ are possible only if $\Box_{\tilde{T}}\Phi^{I}\neq 0$. Following [22], the latter is naturally achieved by introducing some time-dependence to the internal generators $\Phi_{I}$, corresponding to a non-trivial $R$ charge. This is achieved by an ansatz of the form
@@
where the ${\mathcal{K}}_{J}$ are fixed matrices. This respects the full $E(3)$ symmetry of the cosmology, and it allows to stabilize a large hierarchy between the cosmic IR scale and a UV scale set by the Kaluza-Klein modes arising on the compact space $\mathcal{K}$} generated by the $\Phi^{j}$, as demonstrated for k=1k=-1 in [22]. However, this will not affect the equations of motion for the background $\tilde{T}^{\mu}$.
To find physically reasonable solutions for $\tilde{T}^{j}$, it thus appears that quantum corrections are required. Indeed, quantum corrections arising e.g. from the one-loop effective action will certainly modify the equations of motion, and a partial analysis of such effects was carried out for the k=1k=-1 case in [5, 22]. Using covariance arguments, one may reason along the lines of [22] that the one-loop quantum corrections modify the classical equations of (5.5), so that the right-hand side is not vanishing but dependent on the dilaton and the scale $m_{{\mathcal{K}}}^{2}$ of the extra dimension, yielding in particular
@@
Let us also mention that the one-loop effective action on covariant spacetimes arising from the IKKT model was computed in [33] and worked out in some detail for cosmological k=1k=-1 quantum spacetimes in [5] under the assumption that the full ten-dimensional background has a product structure ${\mathcal{M}}^{1,3}\times{\mathcal{K}}$. For related works, see [24, 34, 4, 3].
@@ 6 Discussion
An interesting result is the existence of a background $\beta=c+\tfrac{d}{y^{0}}$ in the covariant gauge, where the higher-spin contributions to the effective metric can be eliminated globally by going to local normal coordinates. This background, which corresponds to an expanding FLRW spacetime with scale factor $a(t)\sim t^{\frac{1}{3}}$, is a sum of the reference background investigated in [12] and the background $\beta_{0}\sim\tfrac{1}{y^{0}}$, which on its own yields a degenerate frame and (inverse) metric.
@@
Finally in Section 5, we have outlined how the cosmological $k=0$ quantum spacetimes under consideration can be embedded into the full ten-dimensional IKKT model. The details of such an embedding are beyond the scope of the present paper. Although we expect that this works similar to the better-known case of k=1k=-1 spacetimes [3, 5, 4, 22], there are some important differences. In particular, the fact that the $k=0$ case is "more commutative" means that time-dependent extra-dimensions will not affect the classical equations of motion of spacetime, so that quantum effects are required to stabilize spacetime.
@@
$T^{\mu}_{k=-1}:=\frac{1}{R}M^{\mu 4},$ (A.1)
which can be generalized and made dynamical by replacing $T^{\mu}_{k=-1}$ by $\tilde{T}^{\mu}_{k=-1}:=f(X^{4})T^{\mu}_{k=-1}$. Here, $f$ is a (well-behaved) function of $X^{4}$, which can be expressed in terms of the cosmological time variable only [27, 28, 22, 34]. The $X^{\mu}$ and $T^{\mu}_{k=-1}$ satisfy the commutation relations
@@
$[T^{\mu}_{k=-1},X^{\nu}]$ $=\frac{\textup{i}X^{4}}{R}\eta^{\mu\nu}.$ (A.2)
The generators $M^{ab}$ and the matrices $X^{a}$ satisfy a number of $SO(4,1)$-invariant constraints in the doubleton representations. Among these, we will make use of the identity (2.3) and [34]
$\varepsilon_{abcde}M^{ab}M^{cd}=\frac{4n}{r}X_{e},\qquad a,b,c,d,e=0,...,4.$ (A.3)
@@
$\{x^{4},x^{\mu}\}=r^{2}Rt_{k=-1}^{\mu}\qquad\text{and}\qquad\{x^{4},t_{k=-1}^{\mu}\}=\frac{x^{\mu}}{R}.$ (A.4)
@@
$x^{a}$ $=R\begin{pmatrix}\cosh(\eta)\cosh(\chi)\\ \cosh(\eta)\sinh(\chi)\bm{\Omega}\\ \sinh(\eta)\end{pmatrix}$, (A.5)
@@
Appendix B Semi-classical generators of rotations and boosts
We determine the semi-classical form of the generators of rotations, $M^{ij}\sim m^{ij}$ and of the boosts, $M^{i0}\sim m^{i0}$ as functions of the vectors $\bm{y}$ and $\bm{t}$ and the $SO(3)$-invariant variables $y^{0}$, $t^{0}$ and $y:=\sqrt{\delta_{ij}y^{i}y^{j}}$, assuming that the parameter $n$ parametrizing the representation is either large, $n\gg 0$, or vanishes $n=0$. Starting point are the concrete forms $m^{\mu\nu}_{k=-1}$ of the semi-classical generators of $SO(3,1)$ in the k=1k=-1 case, which is in the large-$n$-limit given by [34, Eq. (5.4.13)]
$m^{\mu\nu}=-\frac{R}{R^{2}+{x^{4}}^{2}}\Big(x^{4}(x^{\mu}t_{k=-1}^{\nu}-x^{\nu}t_{k=-1}^{\mu})+R\varepsilon^{\mu\nu\alpha\beta}x_{\alpha}t_{k=-1,\beta}\Big),\qquad n\gg 0,$ (B.1)
@@ Appendix B Semi-classical generators of rotations and boosts
In the case $n=0$, we have the simpler form [23]
$m^{\mu\nu}$ $=\frac{R}{x^{4}}(t_{k=-1}^{\mu}x^{\nu}-t_{k=-1}^{\nu}x^{\mu}).$ (B.2)
@@
$\qquad\text{or}\qquad A$ $=\bm{t}\wedge\bm{y}=\ast(\bm{t}\times\bm{y}),\qquad B=\ast\bm{y},\qquad\text{and}\qquad C=\ast\bm{t}.$ (B.3)
@@
$D^{ij}$ $:=A_{kl}(\varepsilon^{ikl}y^{j}-\varepsilon^{jkl}y^{i})\qquad\text{and}\qquad F^{ij}:=A_{kl}(\varepsilon^{ikl}t^{j}-\varepsilon^{jkl}t^{i})$ (B.4)
@@
$\frac{1}{2}F^{ij}$ $=\frac{{y^{0}}^{2}}{r^{2}R^{2}}B^{ij}-y^{0}t^{0}C^{ij}.$ (B.5)
Remark B.1.
In the case of a k=1k=-1 spacetime, where we look for antisymmetric rank two tensors in four spacetime dimensions, invariant under $SO(3,1)$, we only have two independent candidates $A_{4}=t\wedge y$ and $B_{4}=\ast(t\wedge y)$, while in the case of $k=0$, three spatial dimensions and $SO(3)$, we have the three candidates (B.3).
@@
Starting from the form (B.2), the formula (3.19) was proven in [12] for the minimal case $n=0$.
@@
Substituting $y^{\mu}$ and $t^{\mu}$ for $x^{\mu}$ and $t_{k=-1}^{\mu}$ into (B.1) yields [^7] Note that one needs to be careful raising and lowering indices: The indices of $y^{\mu}$ and $x^{\mu}$ are lowered with a different Minkowski metric. In the following, all indices are raised and lowered with the Minkowski metric corresponding to $x^{\mu}$. This does not matter for the spatial components, but we make sure that $x^{0}$, $t^{0}$ and $y^{0}$ do only appear with upper indices in this derivation.
@@
$\qquad+\frac{x^{4}}{R^{2}+{x^{4}}^{2}}(y^{i}m^{j0}-y^{j}m^{i0})-\frac{Rx^{0}}{R^{2}+{x^{4}}^{2}}\varepsilon^{ij}_{\phantom{ij}k}m^{k0}.$ (B.6)
@@
$=\frac{1}{R^{2}+x^{4}y^{0}}\Big(Rx^{4}(x^{0}t^{i}-t^{0}y^{i})+R^{2}\varepsilon^{ijk}y_{j}t_{k}-R\delta^{il}\varepsilon_{ljk}y^{j}m^{k0}\Big).$ (B.7)
@@
$a=-\frac{Rt^{0}}{y^{0}},\qquad b=R\frac{{y^{0}}^{2}+y^{2}-R^{2}}{2{y^{0}}^{2}},\qquad\text{and}\qquad c=\frac{R^{2}}{{y^{0}}^{2}},$ (B.10)

### v2

@@ 3.2 Explicit form of the $SO(3,1)$ generators
It was shown in [Gass:2025tae] that in the minimal representation $n=0$, the generators of rotations $M^{ij}$ are given by [^4] Note that the generators $T^{\mu}$ are no longer hermitian for $n=0$ because $R=\textup{i}r$ in this case. The $SO(3,1)$ generators, however, remain hermitian. $T^{\mu}$ can be made hermitian for $n=0$ via the redefinition $T^{\mu}=\frac{1}{r}(M^{\mu 0}+M^{\mu 4})$, and adjusting the formulas for the $SO(3,1)$ generators accordingly.
@@
where $c$, $d$ and $f$ are functions of the $SO(3)$-invariant quantities $Y^{0}$, $T^{0}$ and $Y^{k}Y_{k}$. However, their semi-classical limits $m^{ij}\sim M^{ij}$ for $n\gg 0$ and also in the minimal case $n=0$ can be determined explicitly. In Appendix C, we derive
@@
A simple explicit realization of the minimal $n=0$ representation in terms of three canonical generators is given in [Ho:2025htr].
@@ 3.3 Fiber projection
Fiber projections of polynomials of higher order in $t^{\mu}$ can be computed using Wick type formulas as in [Steinacker:2024unq, Steinacker:2019awe].
@@ 3.4 Derivatives
In this subsection, we will construct suitable derivative operators acting on the algebra of $\mathfrak{hs}$-valued functions on spacetime, analogous to the construction for k=1k=-1 in [Steinacker:2024unq]. We will always assume $n\gg 0$. To obtain more transparent formulas, it is sometimes convenient to use the variables $x^{a}$ suitable to describe the fuzzy hyperboloid and k=1k=-1 spacetimes instead of $y^{\mu}$. However, in all formulas of the following considerations, $x^{a}$ should be understood as $x^{a}(y^{\mu})$ via the relations
@@
Moreover, Eq. (3.24) and the Poisson brackets (B.4) yield
@@ 3.5 Reconstruction of divergence-free vector fields
where $\eth_{a}$ is the tangential derivative on $H^{4}$ defined in (2.7); and that conversely any vector field $V^{\mu}$ on ${\mathcal{M}}^{1,3}_{k=-1}$ possesses a lift to $H^{4}$ so that (3.30) is satisfied [Steinacker:2024unq]. Note that the vector fields may be $\mathfrak{hs}$ valued. Therefore, there is a correspondence between tangential and divergence-free vector fields on $H^{4}$ to divergence-free vector fields on ${\mathcal{M}}^{1,3}_{k=-1}$. In the following, we establish the analogous correspondence for the $k=0$ case.
@@ 4.5 Local normal coordinates
Due to the $\mathfrak{hs}$ components, the physical meaning of a metric or frame is not clear a priori. As shown in [Steinacker:2024unq], it is always possible to locally eliminate these $\mathfrak{hs}$ components, in local normal coordinates. We illustrate this in the following for the present backgrounds.
@@ 4.7 Physical scales and consistency conditions
On the non-commutative level, the leading contribution to $\lvert[Y^{\mu},Y^{\nu}][Y_{\mu},Y_{\nu}]\rvert$ is the mixed spatial-temporal entry [Steinacker:2024unq]. This should remain true on the semi-classical level, which is easy to verify if the ratio $\tfrac{\alpha}{\beta}$ does not grow too fast. In this case, we generically find
@@
Finally, we require the dilaton to be either constant or increasing with $y^{0}$, corresponding to weak Yang-Mills coupling $g_{\rm YM}\sim\rho^{-1}$ [Steinacker:2024huv]. Using the formula (4.13), this means that $3(1+\mu)+\lambda\geq 0$. This condition might not be strictly necessary, but we will find later that it only excludes some shrinking FLRW spacetimes, which are anyway not interesting from a physical point of view.
@@ 5 FLRW quantum spacetimes and the IKKT model
Let us finally describe how the cosmological quantum spacetimes described above could be incorporated into the IKKT model [^6] It is sometimes questioned whether the IKKT mode admits any non-trivial backgrounds or saddle points with reduced dimension. These doubts are resolved by the example of flat $D3$ BPS branes realized by the Moyal-Weyl quantum plane ${\mathbb{R}}^{3,1}_{\theta}$, leading to noncommutative ${\mathcal{N}}=4$ SYM [Aoki:1999vr].. To this end, consider the bosonic part of the IKKT action (1.5),
@@
Since $\Box_{\Phi}=[\Phi^{I},[\Phi_{I},\cdot]]$ is a positive operator, non-trivial solutions for the $\Phi^{I}$ are possible only if $\Box_{\tilde{T}}\Phi^{I}\neq 0$. Following [MantaStein_deformed], the latter is naturally achieved by introducing some time-dependence to the internal generators $\Phi_{I}$, corresponding to a non-trivial $R$ charge. This is achieved by an ansatz of the form
@@
where the ${\mathcal{K}}_{J}$ are fixed matrices. This respects the full $E(3)$ symmetry of the cosmology, and it allows to stabilize a large hierarchy between the cosmic IR scale and a UV scale set by the Kaluza-Klein modes arising on the compact space $\mathcal{K}$} generated by the $\Phi^{j}$, as demonstrated for k=1k=-1 in [MantaStein_deformed]. However, this will not affect the equations of motion for the background $\tilde{T}^{\mu}$.
To find physically reasonable solutions for $\tilde{T}^{j}$, it thus appears that quantum corrections are required. Indeed, quantum corrections arising e.g. from the one-loop effective action will certainly modify the equations of motion, and a partial analysis of such effects was carried out for the k=1k=-1 case in [Battista:2023glw, MantaStein_deformed]. Using covariance arguments, one may reason along the lines of [MantaStein_deformed] that the one-loop quantum corrections modify the classical equations of (5.5), so that the right-hand side is not vanishing but dependent on the dilaton and the scale $m_{{\mathcal{K}}}^{2}$ of the extra dimension, yielding in particular
@@
Let us also mention that the one-loop effective action on covariant spacetimes arising from the IKKT model was computed in [Steinacker:2023myp] and worked out in some detail for cosmological k=1k=-1 quantum spacetimes in [Battista:2023glw] under the assumption that the full ten-dimensional background has a product structure ${\mathcal{M}}^{1,3}\times{\mathcal{K}}$. For related works, see [Manta:2024vol, Steinacker:2024unq, Battista:2022hqn, Battista:2022vvl].
@@ 6 Discussion
An interesting result is the existence of a background $\beta=c+\tfrac{d}{y^{0}}$ in the covariant gauge, where the higher-spin contributions to the effective metric can be eliminated globally by going to local normal coordinates. This background, which corresponds to an expanding FLRW spacetime with scale factor $a(t)\sim t^{\frac{1}{3}}$, is a sum of the reference background investigated in [Gass:2025tae] and the background $\beta_{0}\sim\tfrac{1}{y^{0}}$, which on its own yields a degenerate frame and (inverse) metric.
@@
Finally in Section 5, we have outlined how the cosmological $k=0$ quantum spacetimes under consideration can be embedded into the full ten-dimensional IKKT model. The details of such an embedding are beyond the scope of the present paper. Although we expect that this works similar to the better-known case of k=1k=-1 spacetimes [Battista:2022vvl, Battista:2023glw, Battista:2022hqn, MantaStein_deformed], there are some important differences. In particular, the fact that the $k=0$ case is "more commutative" means that time-dependent extra-dimensions will not affect the classical equations of motion of spacetime, so that quantum effects are required to stabilize spacetime.
Appendix A Remarks on the quantization and the semi-classical correspondence
We discuss the explicit quantization and the semi-classical correspondence for the symplectic space ${\mathbb{C}}P^{1,2}$ underlying the covariant quantum spacetime ${\mathcal{M}}^{1,3}$.
Since ${\mathbb{C}}P^{1,2}$ is a coadjoint orbit of $SU(2,2)$, it carries a natural (Kirillov-Kostant-Souriau) invariant symplectic form. One then expects that a $SU(2,2)$-equivariant quantization in terms of operators on suitable representations exists. The appropriate representations are easily recognized as the doubleton representations, which can be constructed conveniently using an oscillator construction; this is discussed in detail in [Manta:2025inq, Steinacker:2024unq]. This provides a quantization map where polynomial functions on ${\mathbb{C}}P^{1,2}$ generated by $m^{ab}\in\mathfrak{su}(2,2)^{*}$ are mapped to (symmetrized) polynomials in the generators $M^{ab}$ of $\mathfrak{su}(2,2)$ in the chosen representation, respecting the Poincare-Hilbert series i.e. the number of independent polynomials of any given degree. Using the oscillator construction and its Poisson version, it is then straightforward to verify (1.3).
Another – often more useful – way to define quantization is based on (quasi-)coherent i.e. optimally localized states $|x\rangle$ via
$\Phi=\int\Omega\phi(x)|x\rangle\langle x|,\hskip 18.49988pt\phi(x)\approx\langle x|\Phi|x\rangle\ .$ (A.1)
For compact coadjoint orbits of semi-simple Lie groups, such coherent states are obtained as group orbits of the highest weight state of some representation [Perelomov:1986uhd]. This orbit is then isomorphic to the corresponding coadjoint orbit. For non-compact groups, the story is more complicated, and there is no obvious choice of coherent states. In the present case there appears to be a natural candidate given by the lowest weight state of the doubleton representations of $SU(2,2)$; however their group orbit is not isomorphic to ${\mathbb{C}}P^{1,2}$ but higher-dimensional, and moreover these states are not optimally localized. Nevertheless, it is possible to construct quasi-coherent states which are optimally localized in some local patch of ${\mathbb{C}}P^{1,2}$, i.e. saturating the uncertainty implied by the symplectic volume form $\Omega$; we refer to [Manta:2025inq] for details. This allows to use (A.1) in the local patch, which justifies the semi-classical treatment of ${\mathcal{M}}^{1,3}$ in the present paper.
@@ Appendix A Remarks on the quantization and the semi-classical correspondence
$T^{\mu}_{k=-1}:=\frac{1}{R}M^{\mu 4},$ (B.1)
which can be generalized and made dynamical by replacing $T^{\mu}_{k=-1}$ by $\tilde{T}^{\mu}_{k=-1}:=f(X^{4})T^{\mu}_{k=-1}$. Here, $f$ is a (well-behaved) function of $X^{4}$, which can be expressed in terms of the cosmological time variable only [Sperling:2019xar, Sperling:2018xrm, MantaStein_deformed, Steinacker:2024unq]. The $X^{\mu}$ and $T^{\mu}_{k=-1}$ satisfy the commutation relations
@@
$[T^{\mu}_{k=-1},X^{\nu}]$ $=\frac{\textup{i}X^{4}}{R}\eta^{\mu\nu}.$ (B.2)
The generators $M^{ab}$ and the matrices $X^{a}$ satisfy a number of $SO(4,1)$-invariant constraints in the doubleton representations. Among these, we will make use of the identity (2.3) and [Steinacker:2024unq]
$\varepsilon_{abcde}M^{ab}M^{cd}=\frac{4n}{r}X_{e},\qquad a,b,c,d,e=0,...,4.$ (B.3)
@@
$\{x^{4},x^{\mu}\}=r^{2}Rt_{k=-1}^{\mu}\qquad\text{and}\qquad\{x^{4},t_{k=-1}^{\mu}\}=\frac{x^{\mu}}{R}.$ (B.4)
@@
$x^{a}$ $=R\begin{pmatrix}\cosh(\eta)\cosh(\chi)\\ \cosh(\eta)\sinh(\chi)\bm{\Omega}\\ \sinh(\eta)\end{pmatrix}$, (B.5)
@@
Appendix C Semi-classical generators of rotations and boosts
We determine the semi-classical form of the generators of rotations, $M^{ij}\sim m^{ij}$ and of the boosts, $M^{i0}\sim m^{i0}$ as functions of the vectors $\bm{y}$ and $\bm{t}$ and the $SO(3)$-invariant variables $y^{0}$, $t^{0}$ and $y:=\sqrt{\delta_{ij}y^{i}y^{j}}$, assuming that the parameter $n$ parametrizing the representation is either large, $n\gg 0$, or vanishes $n=0$. Starting point are the concrete forms $m^{\mu\nu}_{k=-1}$ of the semi-classical generators of $SO(3,1)$ in the k=1k=-1 case, which is in the large-$n$-limit given by [Steinacker:2024unq, Eq. (5.4.13)]
$m^{\mu\nu}=-\frac{R}{R^{2}+{x^{4}}^{2}}\Big(x^{4}(x^{\mu}t_{k=-1}^{\nu}-x^{\nu}t_{k=-1}^{\mu})+R\varepsilon^{\mu\nu\alpha\beta}x_{\alpha}t_{k=-1,\beta}\Big),\qquad n\gg 0,$ (C.1)
@@ Appendix C Semi-classical generators of rotations and boosts
In the case $n=0$, we have the simpler form [Manta:2025inq]
$m^{\mu\nu}$ $=\frac{R}{x^{4}}(t_{k=-1}^{\mu}x^{\nu}-t_{k=-1}^{\nu}x^{\mu}).$ (C.2)
@@
$\qquad\text{or}\qquad A$ $=\bm{t}\wedge\bm{y}=\ast(\bm{t}\times\bm{y}),\qquad B=\ast\bm{y},\qquad\text{and}\qquad C=\ast\bm{t}.$ (C.3)
@@
$D^{ij}$ $:=A_{kl}(\varepsilon^{ikl}y^{j}-\varepsilon^{jkl}y^{i})\qquad\text{and}\qquad F^{ij}:=A_{kl}(\varepsilon^{ikl}t^{j}-\varepsilon^{jkl}t^{i})$ (C.4)
@@
$\frac{1}{2}F^{ij}$ $=\frac{{y^{0}}^{2}}{r^{2}R^{2}}B^{ij}-y^{0}t^{0}C^{ij}.$ (C.5)
Remark C.1.
In the case of a k=1k=-1 spacetime, where we look for antisymmetric rank two tensors in four spacetime dimensions, invariant under $SO(3,1)$, we only have two independent candidates $A_{4}=t\wedge y$ and $B_{4}=\ast(t\wedge y)$, while in the case of $k=0$, three spatial dimensions and $SO(3)$, we have the three candidates (C.3).
@@
Starting from the form (C.2), the formula (3.19) was proven in [Gass:2025tae] for the minimal case $n=0$.
@@
Substituting $y^{\mu}$ and $t^{\mu}$ for $x^{\mu}$ and $t_{k=-1}^{\mu}$ into (C.1) yields [^7] Note that one needs to be careful raising and lowering indices: The indices of $y^{\mu}$ and $x^{\mu}$ are lowered with a different Minkowski metric. In the following, all indices are raised and lowered with the Minkowski metric corresponding to $x^{\mu}$. This does not matter for the spatial components, but we make sure that $x^{0}$, $t^{0}$ and $y^{0}$ do only appear with upper indices in this derivation.
@@
$\qquad+\frac{x^{4}}{R^{2}+{x^{4}}^{2}}(y^{i}m^{j0}-y^{j}m^{i0})-\frac{Rx^{0}}{R^{2}+{x^{4}}^{2}}\varepsilon^{ij}_{\phantom{ij}k}m^{k0}.$ (C.6)
@@
$=\frac{1}{R^{2}+x^{4}y^{0}}\Big(Rx^{4}(x^{0}t^{i}-t^{0}y^{i})+R^{2}\varepsilon^{ijk}y_{j}t_{k}-R\delta^{il}\varepsilon_{ljk}y^{j}m^{k0}\Big).$ (C.7)
@@
$a=-\frac{Rt^{0}}{y^{0}},\qquad b=R\frac{{y^{0}}^{2}+y^{2}-R^{2}}{2{y^{0}}^{2}},\qquad\text{and}\qquad c=\frac{R^{2}}{{y^{0}}^{2}},$ (C.10)